# Training-Free Bayesianization for Low-Rank Adapters of Large Language Models

### v1

@@ C.2 TFB Beyond Low-Rank Isotropic Gaussians
ACC ($\uparrow$) MLE 68.99$\pm$0.58 69.10$\pm$2.84 85.65$\pm$0.92 74.53$\pm$0.66 81.52$\pm$0.25 86.53$\pm$0.28 66.20$\pm$0.87 75.12$\pm$0.85 40.62$\pm$2.25 28.82$\pm$1.30
 + TFB* (C-STD) 68.43$\pm$0.78 68.13$\pm$0.52 84.45$\pm$0.10 74.10$\pm$1.57 81.87$\pm$0.23 87.49$\pm$0.34 68.69$\pm$0.85 73.59$\pm$1.44 38.67$\pm$2.89 29.33$\pm$4.73
 + TFB* (FR) 67.19$\pm$1.46 69.53$\pm$0.55 85.29$\pm$0.79 73.05$\pm$0.67 81.40$\pm$0.21 86.67$\pm$0.03 68.55$\pm$0.20 72.85$\pm$0.00 44.53$\pm$2.34 35.16$\pm$3.91
 + TFB* (Ours) 66.59$\pm$2.13 65.77$\pm$2.44 85.74$\pm$0.63 75.63$\pm$0.48 83.00$\pm$1.31 87.63$\pm$0.23 69.48$\pm$0.85 75.06$\pm$1.34 38.67$\pm$2.52 29.00$\pm$7.81
MAP 68.62$\pm$0.71 67.59$\pm$0.40 86.55$\pm$0.55 75.61$\pm$0.71 81.38$\pm$0.65 86.50$\pm$0.41 69.59$\pm$0.33 75.47$\pm$0.73 44.79$\pm$0.00 28.47$\pm$1.20
 + TFB* (C-STD) 68.59$\pm$1.10 66.55$\pm$0.59 85.15$\pm$0.41 73.36$\pm$0.44 80.60$\pm$0.69 87.44$\pm$0.49 68.69$\pm$0.85 73.59$\pm$1.44 38.67$\pm$2.89 29.33$\pm$4.73
 + TFB* (FR) 70.12$\pm$0.55 68.73$\pm$0.37 85.31$\pm$0.94 73.01$\pm$0.39 82.28$\pm$1.26 86.52$\pm$0.06 67.77$\pm$0.98 70.90$\pm$1.17 46.88$\pm$0.00 36.72$\pm$0.78
 + TFB* (Ours) 69.12$\pm$1.42 66.89$\pm$2.68 86.27$\pm$0.93 75.16$\pm$0.73 82.53$\pm$1.42 87.60$\pm$0.34 69.59$\pm$1.22 75.29$\pm$0.83 37.67$\pm$1.53 29.00$\pm$6.00
BLoB 69.07$\pm$0.34 68.81$\pm$1.09 85.56$\pm$0.35 73.69$\pm$0.17 81.52$\pm$0.74 86.99$\pm$0.24 67.71$\pm$1.13 76.37$\pm$0.80 44.79$\pm$1.47 31.60$\pm$2.73
BLoB-Mean 70.89$\pm$0.82 70.83$\pm$1.57 86.68$\pm$0.60 74.55$\pm$1.94 82.73$\pm$0.41 86.80$\pm$0.23 69.93$\pm$1.20 76.88$\pm$0.41 41.67$\pm$2.25 31.94$\pm$1.77
 + TFB* (C-STD) 70.06$\pm$0.95 68.36$\pm$1.52 84.98$\pm$0.89 74.45$\pm$0.93 82.87$\pm$0.46 87.42$\pm$0.00 68.02$\pm$0.70 76.70$\pm$0.44 45.33$\pm$1.53 31.00$\pm$3.61
 + TFB* (FR) 68.39$\pm$0.54 67.58$\pm$1.46 86.00$\pm$0.93 73.93$\pm$1.31 82.51$\pm$1.00 86.67$\pm$0.03 69.73$\pm$0.98 74.51$\pm$0.88 45.31$\pm$0.00 38.28$\pm$0.78
 + TFB* (Ours) 68.99$\pm$0.34 69.26$\pm$1.22 86.97$\pm$0.30 75.50$\pm$0.67 82.87$\pm$1.03 87.16$\pm$0.29 70.27$\pm$0.89 77.35$\pm$0.41 37.33$\pm$2.31 30.33$\pm$2.08
ECE ($\downarrow$) MLE 29.83$\pm$0.58 29.00$\pm$1.97 13.12$\pm$1.39 20.62$\pm$0.74 12.55$\pm$0.46 3.18$\pm$0.09 22.20$\pm$0.39 16.47$\pm$0.86 21.72$\pm$0.30 29.60$\pm$1.29
 + TFB* (C-STD) 16.59$\pm$0.31 13.15$\pm$0.90 4.45$\pm$0.31 12.00$\pm$2.15 4.17$\pm$0.99 1.25$\pm$0.30 7.09$\pm$2.38 5.56$\pm$0.24 14.31$\pm$4.25 23.20$\pm$3.44
 + TFB* (FR) 19.55$\pm$0.87 15.26$\pm$1.19 5.98$\pm$0.71 12.96$\pm$0.48 5.30$\pm$1.09 1.21$\pm$0.09 10.41$\pm$0.83 9.20$\pm$0.30 18.50$\pm$4.20 22.35$\pm$3.35
 + TFB* (Ours) 5.62$\pm$1.05 6.63$\pm$0.97 4.96$\pm$0.83 4.77$\pm$0.59 3.64$\pm$1.33 1.39$\pm$0.34 7.39$\pm$0.95 5.07$\pm$0.10 12.55$\pm$0.95 19.30$\pm$6.49
MAP 29.76$\pm$0.87 29.42$\pm$0.68 12.07$\pm$0.55 23.07$\pm$0.14 13.26$\pm$0.82 3.16$\pm$0.23 19.31$\pm$1.46 15.68$\pm$0.51 17.55$\pm$1.95 30.25$\pm$2.18
 + TFB* (C-STD) 17.42$\pm$0.94 11.48$\pm$0.71 4.01$\pm$0.97 10.38$\pm$0.40 4.16$\pm$0.59 1.07$\pm$0.35 7.09$\pm$2.38 5.56$\pm$0.24 14.31$\pm$4.25 23.20$\pm$3.44
 + TFB* (FR) 16.36$\pm$2.65 16.24$\pm$1.27 5.97$\pm$1.10 11.89$\pm$3.69 3.77$\pm$0.80 1.16$\pm$0.08 11.08$\pm$0.62 9.72$\pm$0.55 14.71$\pm$1.10 24.27$\pm$2.57
 + TFB* (Ours) 9.89$\pm$0.41 6.45$\pm$1.61 3.61$\pm$1.27 4.27$\pm$0.48 2.96$\pm$0.05 1.19$\pm$0.38 5.01$\pm$1.15 4.33$\pm$0.31 12.08$\pm$0.67 19.53$\pm$5.25
BLoB 9.35$\pm$1.37 9.59$\pm$1.88 3.64$\pm$0.53 3.01$\pm$0.12 3.77$\pm$1.47 1.41$\pm$0.19 9.55$\pm$0.40 5.48$\pm$1.27 9.77$\pm$1.35 18.29$\pm$1.35
BLoB-Mean 20.62$\pm$0.83 20.61$\pm$1.16 9.43$\pm$0.38 11.23$\pm$0.69 8.36$\pm$0.38 2.46$\pm$0.07 14.00$\pm$1.02 10.70$\pm$0.39 15.05$\pm$0.77 22.90$\pm$2.27
 + TFB* (C-STD) 15.53$\pm$1.51 6.49$\pm$1.18 3.94$\pm$0.13 5.94$\pm$0.65 3.76$\pm$0.97 3.51$\pm$0.00 9.79$\pm$1.00 5.11$\pm$0.83 8.67$\pm$1.91 18.30$\pm$0.53
 + TFB* (FR) 12.31$\pm$0.70 7.77$\pm$1.11 4.45$\pm$0.80 7.17$\pm$1.18 4.29$\pm$0.89 1.21$\pm$0.09 6.91$\pm$0.62 6.04$\pm$0.47 10.12$\pm$3.06 14.05$\pm$0.98
 + TFB* (Ours) 4.15$\pm$0.67 4.77$\pm$1.58 2.57$\pm$0.56 4.03$\pm$1.14 3.50$\pm$0.53 1.41$\pm$0.68 8.35$\pm$0.51 5.39$\pm$0.21 14.01$\pm$4.76 17.10$\pm$1.23
NLL ($\downarrow$) MLE 3.17$\pm$0.37 2.85$\pm$0.27 1.17$\pm$0.13 0.95$\pm$0.07 0.73$\pm$0.03 0.32$\pm$0.00 1.16$\pm$0.00 0.92$\pm$0.03 1.56$\pm$0.06 1.66$\pm$0.05
 + TFB* (C-STD) 1.15$\pm$0.19 1.26$\pm$0.06 0.54$\pm$0.06 0.67$\pm$0.05 0.55$\pm$0.01 0.30$\pm$0.00 0.87$\pm$0.05 0.72$\pm$0.01 1.38$\pm$0.03 1.49$\pm$0.03
 + TFB* (FR) 1.25$\pm$0.15 1.32$\pm$0.11 0.54$\pm$0.05 0.69$\pm$0.01 0.55$\pm$0.01 0.31$\pm$0.00 0.91$\pm$0.02 0.77$\pm$0.01 1.34$\pm$0.02 1.50$\pm$0.04
 + TFB* (Ours) 0.67$\pm$0.00 1.05$\pm$0.06 0.52$\pm$0.01 0.56$\pm$0.01 0.56$\pm$0.02 0.30$\pm$0.00 0.85$\pm$0.03 0.73$\pm$0.04 1.41$\pm$0.04 1.49$\pm$0.02
MAP 2.46$\pm$0.34 2.66$\pm$0.11 0.90$\pm$0.05 1.62$\pm$0.29 0.75$\pm$0.01 0.33$\pm$0.00 1.10$\pm$0.07 0.93$\pm$0.04 1.55$\pm$0.06 1.65$\pm$0.03
 + TFB* (C-STD) 1.07$\pm$0.04 1.20$\pm$0.05 0.53$\pm$0.03 0.62$\pm$0.01 0.56$\pm$0.02 0.30$\pm$0.01 0.87$\pm$0.05 0.72$\pm$0.01 1.38$\pm$0.03 1.49$\pm$0.03
 + TFB* (FR) 0.89$\pm$0.20 1.15$\pm$0.23 0.50$\pm$0.08 0.68$\pm$0.11 0.52$\pm$0.03 0.31$\pm$0.00 0.93$\pm$0.00 0.76$\pm$0.00 1.40$\pm$0.02 1.51$\pm$0.03
 + TFB* (Ours) 0.72$\pm$0.02 1.04$\pm$0.07 0.51$\pm$0.01 0.56$\pm$0.01 0.56$\pm$0.02 0.30$\pm$0.00 0.83$\pm$0.03 0.73$\pm$0.04 1.41$\pm$0.04 1.49$\pm$0.03
BLoB 0.63$\pm$0.01 0.78$\pm$0.02 0.40$\pm$0.01 0.54$\pm$0.00 0.50$\pm$0.01 0.31$\pm$0.00 0.83$\pm$0.01 0.60$\pm$0.01 1.38$\pm$0.01 1.46$\pm$0.02
BLoB-Mean 0.91$\pm$0.10 1.19$\pm$0.02 0.56$\pm$0.01 0.60$\pm$0.01 0.56$\pm$0.02 0.32$\pm$0.00 0.89$\pm$0.02 0.67$\pm$0.02 1.44$\pm$0.00 1.53$\pm$0.02
 + TFB* (C-STD) 0.70$\pm$0.04 0.83$\pm$0.02 0.43$\pm$0.01 0.53$\pm$0.01 0.47$\pm$0.01 0.31$\pm$0.00 0.79$\pm$0.02 0.63$\pm$0.02 1.31$\pm$0.03 1.44$\pm$0.03
 + TFB* (FR) 0.67$\pm$0.01 0.83$\pm$0.01 0.41$\pm$0.01 0.55$\pm$0.02 0.48$\pm$0.01 0.31$\pm$0.00 0.79$\pm$0.00 0.61$\pm$0.02 1.35$\pm$0.03 1.46$\pm$0.05
 + TFB* (Ours) 0.61$\pm$0.01 0.88$\pm$0.02 0.41$\pm$0.01 0.54$\pm$0.01 0.48$\pm$0.01 0.30$\pm$0.00 0.80$\pm$0.02 0.64$\pm$0.02 1.37$\pm$0.03 1.45$\pm$0.06
C.3 TFB Beyond the Llama2-7B Backbone
@@
Table 10: TFB Performances with various LLM backbones (Touvron et al., 2023a, b; Dubey et al., 2024; Jiang et al., 2023), where Accuracy (ACC) and Expected Calibration Error (ECE) are reported in percentages. The MLE training for each different backbone is conducted for 2 epochs on the concatenated dataset of six commonsense reasoning tasks, with a shared hyperparameter setting; “Fewer Epochs” represents training for 1 epoch. We sample $N=10$ for TFB during inference and rows with shading indicate training-free Bayesianization methods that use a pre-trained LoRA as their mean. $\uparrow$” and $\downarrow$” indicate that higher and lower values are preferred, respectively. Boldface and underlining denote the best and the second-best performance, respectively.
@@
The results demonstrate that TFB consistently improves model calibration across different backbones while maintaining competitive accuracy. Specifically, for Llama2-7B, TFB reduces the ECE from 4.50 to 1.24 on the combined dataset while preserving the accuracy (81.32 vs 81.41). Similar improvements are observed with Llama3-8B, Llama3.1-8B, and Mistral-7B-v0.3, where TFB achieves better calibration than both the full training and early stopping baselines without sacrificing performance, suggesting its effectiveness as a general approach for enhancing LLM calibration.
C.4 TFB Beyond the Naive LoRA
  • VeRA (Kopiczko et al., 2023): Uses shared low-rank matrices ${\bm{B}}$ and ${\bm{A}}$ across layers, with layer-specific trainable scalar vector ${\bm{d}}$ and bias vector ${\bm{b}}$. Concretely, the parameterization of VeRA’s updated weight matrix ${\bm{W}}$ is modeled as:
    Hence after the fine-tuning of VeRA, we can easily regroup the weight matrices into $\{{\bm{B}}^{\prime}=[\operatorname{diag}({\bm{b}})]{\bm{B}}[\operatorname{diag}({\bm{d}})],{\bm{A}}^{\prime}={\bm{A}}\}$, and apply the TFB Bayesianization scheme illustrated in Algorithm 1.
  • PiSSA (Meng et al., 2024): Employs an alternative initialization scheme while maintaining LoRA’s parameterization and training procedure. Hence the TFB process for PiSSA is trivial.
The results in Table 11 show that TFB consistently improves calibration across different LoRA variants while preserving model performance. Notably, when applied to the standard LoRA, TFB significantly reduces the ECE from 4.74 to 1.05 on the combined dataset with minimal impact on accuracy (86.45 vs 86.70). Similar improvements are observed with VeRA and PiSSA variants, where TFB achieves better calibration (reducing ECE to 1.44 and 1.17 respectively) while maintaining comparable accuracy levels. These results demonstrate that TFB can effectively enhance model calibration across different LoRA architectures without compromising their performance.
Table 11: Performance of TFB when applied to variants of LoRAs (Hu et al., 2022; Zhang et al., 2023; Kopiczko et al., 2023; Meng et al., 2024), where Accuracy (ACC) and Expected Calibration Error (ECE) are reported in percentages. The MLE training for each LoRA variant is conducted with pre-trained Llama3.1-8B model for 2 epochs on the concatenated dataset of six commonsense reasoning tasks, with a shared hyperparameter setting. We sample $N=10$ for TFB during inference and rows with shading indicate training-free Bayesianization methods that use a pre-trained LoRA as their mean. $\uparrow$” and $\downarrow$” indicate that higher and lower values are preferred, respectively. Boldface denotes the best performance.
@@ C.4 TFB Beyond the Naive LoRA
C.5 Improving the Efficiency of TFB
Table 12: Performance of Last-Layer TFB (LL TFB) applied to LoRA on Llama2-7B pre-trained weights, where Accuracy (ACC) and Expected Calibration Error (ECE) are reported in percentages. The evaluation is done across six common-sense reasoning tasks with a shared hyper-parameter setting after 5,000 gradient steps. We sample $N$ times during inference in the sampling-based methods. Rows with shading indicate training-free Bayesianization methods that use a pre-trained LoRA as their mean. $\uparrow$” and $\downarrow$” indicate that higher and lower values are preferred, respectively. Boldface and underlining denote the best and the second-best performance, respectively.

### v2

@@ F.10 Additional Results on Llama2-7B
ACC ($\uparrow$) MCD 78.03$\pm$0.61 81.64$\pm$1.79 91.37$\pm$0.38 83.18$\pm$0.84 87.20$\pm$1.02 89.93$\pm$0.16 81.42$\pm$1.38 87.27$\pm$0.84 47.92$\pm$2.25 46.53$\pm$0.49
ENS 78.82$\pm$0.52 82.55$\pm$0.42 91.84$\pm$0.36 83.99$\pm$0.74 87.37$\pm$0.67 90.50$\pm$0.14 79.62$\pm$0.57 86.56$\pm$0.60 49.65$\pm$3.22 44.44$\pm$1.96
LAP BP 76.05$\pm$0.92 79.95$\pm$0.42 90.73$\pm$0.08 82.83$\pm$0.85 87.90$\pm$0.20 89.36$\pm$0.52 81.08$\pm$1.20 87.21$\pm$1.20 48.26$\pm$3.93 46.18$\pm$1.30
MLE - 77.87$\pm$0.54 81.08$\pm$0.48 91.67$\pm$0.36 82.30$\pm$0.53 87.90$\pm$0.87 89.58$\pm$0.26 81.48$\pm$2.41 86.83$\pm$0.87 45.83$\pm$0.85 42.36$\pm$1.77
 + TFB (${\mathcal{D}}_{\text{train}}$) 77.44$\pm$0.30 82.53$\pm$1.00 91.33$\pm$0.37 82.53$\pm$0.56 88.53$\pm$0.57 89.75$\pm$0.25 79.76$\pm$1.24 85.52$\pm$0.56 44.33$\pm$4.03 37.00$\pm$2.16
 + TFB (${\mathcal{D}}_{\text{test}}$) 77.07$\pm$0.48 82.27$\pm$0.68 91.33$\pm$0.45 82.51$\pm$0.86 88.20$\pm$0.16 89.60$\pm$0.12 81.82$\pm$1.48 86.67$\pm$0.73 47.33$\pm$1.70 40.33$\pm$2.62
MAP - 76.90$\pm$0.97 81.08$\pm$2.48 91.61$\pm$0.44 82.59$\pm$0.28 85.73$\pm$0.19 90.09$\pm$0.28 79.98$\pm$0.87 86.58$\pm$0.79 43.40$\pm$4.98 38.54$\pm$3.40
 + TFB (${\mathcal{D}}_{\text{train}}$) 76.43$\pm$0.72 82.80$\pm$1.42 91.39$\pm$0.37 82.64$\pm$0.58 86.00$\pm$0.16 89.96$\pm$0.18 80.61$\pm$1.24 86.30$\pm$0.89 45.33$\pm$2.87 35.67$\pm$4.11
 + TFB (${\mathcal{D}}_{\text{test}}$) 76.21$\pm$0.88 82.40$\pm$1.70 91.64$\pm$0.74 82.29$\pm$0.85 86.07$\pm$0.19 89.99$\pm$0.13 80.85$\pm$1.12 86.30$\pm$0.89 44.67$\pm$1.70 36.00$\pm$4.32
BLoB 76.45$\pm$0.37 82.32$\pm$1.15 91.14$\pm$0.54 82.01$\pm$0.56 87.57$\pm$0.21 89.65$\pm$0.15 79.75$\pm$0.43 87.13$\pm$0.00 42.71$\pm$3.71 44.79$\pm$6.64
BLoB-Mean 77.72$\pm$0.12 82.60$\pm$0.60 91.64$\pm$0.55 83.92$\pm$0.48 88.00$\pm$0.80 89.86$\pm$0.05 82.06$\pm$1.15 88.54$\pm$0.31 39.93$\pm$5.20 39.93$\pm$4.02
 + TFB (${\mathcal{D}}_{\text{train}}$) 77.81$\pm$0.36 83.33$\pm$0.19 91.76$\pm$0.48 83.81$\pm$0.39 87.80$\pm$0.16 90.11$\pm$0.28 82.93$\pm$1.54 87.64$\pm$0.51 39.67$\pm$7.32 37.33$\pm$6.65
 + TFB (${\mathcal{D}}_{\text{test}}$) 77.65$\pm$0.65 83.20$\pm$1.18 91.94$\pm$0.09 83.41$\pm$0.62 87.73$\pm$0.38 89.71$\pm$0.06 82.93$\pm$1.47 87.15$\pm$0.34 37.67$\pm$4.50 37.00$\pm$6.48
ECE ($\downarrow$) MCD 16.13$\pm$0.54 13.69$\pm$1.11 6.73$\pm$0.71 13.05$\pm$0.99 9.76$\pm$0.71 7.95$\pm$0.17 13.63$\pm$1.18 9.27$\pm$0.60 30.91$\pm$3.57 33.08$\pm$1.40
ENS 14.72$\pm$0.17 13.45$\pm$1.19 6.59$\pm$0.45 11.17$\pm$0.92 8.17$\pm$0.86 7.35$\pm$0.55 11.37$\pm$1.82 7.21$\pm$1.13 18.92$\pm$6.03 26.80$\pm$3.23
LAP BP 4.18$\pm$0.11 9.26$\pm$3.08 5.27$\pm$0.51 3.50$\pm$0.78 8.93$\pm$0.34 1.93$\pm$0.22 7.83$\pm$1.49 7.80$\pm$1.99 14.49$\pm$0.57 13.17$\pm$2.14
MLE - 17.02$\pm$0.46 16.35$\pm$0.68 7.00$\pm$0.53 13.83$\pm$0.65 9.77$\pm$0.81 8.69$\pm$0.21 14.45$\pm$2.19 10.78$\pm$0.50 32.46$\pm$2.60 38.41$\pm$4.44
 + TFB (${\mathcal{D}}_{\text{train}}$) 12.98$\pm$0.37 11.63$\pm$0.68 5.14$\pm$0.14 10.01$\pm$0.70 7.20$\pm$0.47 7.39$\pm$0.26 6.54$\pm$0.53 5.69$\pm$1.64 14.63$\pm$1.46 19.68$\pm$3.27
 + TFB (${\mathcal{D}}_{\text{test}}$) 11.22$\pm$1.15 13.85$\pm$0.66 5.62$\pm$0.86 9.20$\pm$1.41 7.40$\pm$0.60 6.05$\pm$0.44 13.41$\pm$0.72 8.98$\pm$0.51 29.72$\pm$3.01 36.22$\pm$3.50
MAP - 18.71$\pm$0.74 15.77$\pm$1.60 6.62$\pm$0.64 14.26$\pm$0.92 12.19$\pm$0.55 8.40$\pm$0.25 16.46$\pm$0.44 11.36$\pm$0.58 34.79$\pm$3.76 38.50$\pm$2.18
 + TFB (${\mathcal{D}}_{\text{train}}$) 14.95$\pm$0.65 11.27$\pm$2.53 5.76$\pm$0.63 10.97$\pm$1.19 9.70$\pm$0.69 6.86$\pm$0.31 13.25$\pm$0.95 9.22$\pm$0.91 27.21$\pm$2.62 35.91$\pm$4.12
 + TFB (${\mathcal{D}}_{\text{test}}$) 13.08$\pm$1.55 12.85$\pm$2.09 5.03$\pm$0.44 9.28$\pm$1.13 9.86$\pm$0.94 6.44$\pm$0.37 11.07$\pm$1.54 8.02$\pm$1.00 24.51$\pm$3.76 33.25$\pm$5.63
BLoB 9.93$\pm$0.22 5.41$\pm$1.17 2.70$\pm$0.87 4.28$\pm$0.64 2.91$\pm$0.92 2.58$\pm$0.25 5.61$\pm$0.40 2.48$\pm$0.43 16.67$\pm$0.87 12.78$\pm$4.18
BLoB-Mean 15.43$\pm$0.15 12.41$\pm$1.52 4.91$\pm$0.28 9.37$\pm$1.33 6.44$\pm$0.15 6.26$\pm$0.29 11.22$\pm$0.38 6.34$\pm$0.71 26.65$\pm$3.06 25.40$\pm$5.40
 + TFB (${\mathcal{D}}_{\text{train}}$) 8.16$\pm$0.48 6.48$\pm$0.36 2.44$\pm$0.50 3.83$\pm$0.43 2.67$\pm$0.18 3.10$\pm$0.59 6.69$\pm$1.63 3.61$\pm$0.87 18.45$\pm$6.75 20.53$\pm$6.27
 + TFB (${\mathcal{D}}_{\text{test}}$) 7.77$\pm$0.77 6.92$\pm$2.27 2.98$\pm$0.67 2.94$\pm$0.27 3.69$\pm$0.57 1.48$\pm$0.35 4.91$\pm$1.21 3.56$\pm$0.47 17.15$\pm$3.28 16.81$\pm$5.33
NLL ($\downarrow$) MCD 0.83$\pm$0.01 0.99$\pm$0.10 0.45$\pm$0.06 0.64$\pm$0.03 0.62$\pm$0.08 0.49$\pm$0.01 1.03$\pm$0.02 0.61$\pm$0.03 1.91$\pm$0.18 2.02$\pm$0.15
ENS 0.75$\pm$0.02 0.80$\pm$0.11 0.38$\pm$0.03 0.55$\pm$0.02 0.45$\pm$0.05 0.42$\pm$0.05 0.72$\pm$0.07 0.44$\pm$0.03 1.40$\pm$0.18 1.50$\pm$0.13
LAP BP 0.56$\pm$0.00 1.18$\pm$0.02 1.04$\pm$0.01 0.51$\pm$0.00 0.94$\pm$0.00 0.43$\pm$0.00 1.17$\pm$0.01 1.11$\pm$0.00 1.27$\pm$0.01 1.28$\pm$0.00
MLE - 0.88$\pm$0.04 1.20$\pm$0.11 0.46$\pm$0.04 0.68$\pm$0.01 0.61$\pm$0.06 0.52$\pm$0.01 1.07$\pm$0.06 0.72$\pm$0.06 1.91$\pm$0.16 2.25$\pm$0.21
 + TFB (${\mathcal{D}}_{\text{train}}$) 0.68$\pm$0.03 0.85$\pm$0.02 0.33$\pm$0.03 0.53$\pm$0.01 0.46$\pm$0.04 0.42$\pm$0.00 0.66$\pm$0.02 0.44$\pm$0.01 1.39$\pm$0.11 1.49$\pm$0.05
 + TFB (${\mathcal{D}}_{\text{test}}$) 0.63$\pm$0.04 1.00$\pm$0.06 0.36$\pm$0.05 0.51$\pm$0.02 0.46$\pm$0.05 0.36$\pm$0.01 0.83$\pm$0.06 0.51$\pm$0.06 1.59$\pm$0.15 1.94$\pm$0.22
MAP - 0.99$\pm$0.07 1.12$\pm$0.23 0.46$\pm$0.03 0.74$\pm$0.07 0.79$\pm$0.02 0.52$\pm$0.01 1.19$\pm$0.04 0.83$\pm$0.06 1.97$\pm$0.13 2.32$\pm$0.10
 + TFB (${\mathcal{D}}_{\text{train}}$) 0.77$\pm$0.05 0.80$\pm$0.15 0.38$\pm$0.03 0.57$\pm$0.05 0.61$\pm$0.03 0.40$\pm$0.01 0.96$\pm$0.08 0.66$\pm$0.06 1.69$\pm$0.16 2.12$\pm$0.08
 + TFB (${\mathcal{D}}_{\text{test}}$) 0.72$\pm$0.08 0.92$\pm$0.20 0.34$\pm$0.03 0.51$\pm$0.03 0.61$\pm$0.04 0.37$\pm$0.02 0.98$\pm$0.03 0.66$\pm$0.08 1.70$\pm$0.18 2.12$\pm$0.15
BLoB 0.58$\pm$0.00 0.51$\pm$0.03 0.23$\pm$0.01 0.43$\pm$0.01 0.34$\pm$0.01 0.26$\pm$0.01 0.56$\pm$0.02 0.35$\pm$0.02 1.34$\pm$0.04 1.35$\pm$0.10
BLoB-Mean 0.74$\pm$0.02 0.73$\pm$0.04 0.29$\pm$0.03 0.47$\pm$0.03 0.37$\pm$0.02 0.32$\pm$0.02 0.67$\pm$0.07 0.39$\pm$0.03 1.53$\pm$0.13 1.54$\pm$0.15
 + TFB (${\mathcal{D}}_{\text{train}}$) 0.55$\pm$0.01 0.53$\pm$0.04 0.23$\pm$0.02 0.40$\pm$0.01 0.33$\pm$0.02 0.27$\pm$0.01 0.52$\pm$0.05 0.35$\pm$0.02 1.36$\pm$0.13 1.46$\pm$0.11
 + TFB (${\mathcal{D}}_{\text{test}}$) 0.54$\pm$0.01 0.54$\pm$0.04 0.24$\pm$0.01 0.41$\pm$0.01 0.34$\pm$0.01 0.27$\pm$0.01 0.54$\pm$0.03 0.36$\pm$0.02 1.38$\pm$0.10 1.43$\pm$0.10
Table 11: Performance of TFB with accuracy-based evaluation metric ($l$=ACC) using different posterior families applied to the mean of BLoB, based on Llama3.1-8B pre-trained weights. FR: Full-rank isotropic Gaussian noises are applied to $\Delta{\bm{W}}$; C-STD: Standard deviation matrix ${\bm{\Omega}}=[\Omega_{ij}=\sigma_{q}]$ is constant. The evaluation protocol strictly follows Table 1. $\uparrow$” and $\downarrow$” indicate that higher and lower values are preferred, respectively. Boldface and underlining denote the best and the second-best performance, respectively (only for TFB variants).
[…]
Small Shift Large Shift
WG-S ARC-C ARC-E WG-M OBQA BoolQ ARC-C ARC-E Chem Phy
ACC ($\uparrow$) BLoB-Mean 77.72$\pm$0.12 82.60$\pm$0.60 91.64$\pm$0.55 83.92$\pm$0.48 88.00$\pm$0.80 89.86$\pm$0.05 82.06$\pm$1.15 88.54$\pm$0.31 39.93$\pm$5.20 39.93$\pm$4.02
 + TFB (FR) 76.32$\pm$0.45 82.13$\pm$0.38 91.82$\pm$0.65 83.33$\pm$0.85 88.40$\pm$0.16 90.03$\pm$0.18 82.27$\pm$1.05 88.24$\pm$0.37 43.33$\pm$5.79 41.33$\pm$3.30
 + TFB (C-STD) 78.40$\pm$0.36 82.75$\pm$0.16 92.38$\pm$0.08 83.44$\pm$0.46 88.37$\pm$0.50 90.21$\pm$0.04 82.38$\pm$0.89 87.98$\pm$0.30 40.00$\pm$4.68 41.67$\pm$4.12
 + TFB (Final) 75.28$\pm$0.33 82.80$\pm$0.33 91.64$\pm$0.15 81.84$\pm$0.74 88.00$\pm$0.16 89.60$\pm$0.35 82.93$\pm$1.91 86.97$\pm$0.62 36.00$\pm$5.66 36.00$\pm$5.72
ECE ($\downarrow$) BLoB-Mean 15.43$\pm$0.15 12.41$\pm$1.52 4.91$\pm$0.28 9.37$\pm$1.33 6.44$\pm$0.15 6.26$\pm$0.29 11.22$\pm$0.38 6.34$\pm$0.71 26.65$\pm$3.06 25.40$\pm$5.40
 + TFB (FR) 11.49$\pm$0.35 5.74$\pm$1.24 3.21$\pm$0.43 6.43$\pm$0.70 4.22$\pm$0.14 4.90$\pm$0.49 8.85$\pm$1.74 4.18$\pm$0.72 17.63$\pm$4.29 22.31$\pm$5.67
 + TFB (C-STD) 14.07$\pm$0.22 9.85$\pm$1.48 4.17$\pm$0.73 8.94$\pm$0.50 5.48$\pm$0.50 5.72$\pm$0.28 9.29$\pm$0.46 6.13$\pm$0.41 24.78$\pm$3.57 24.51$\pm$2.65
 + TFB (Final) 3.04$\pm$0.12 6.76$\pm$1.47 4.81$\pm$1.18 7.42$\pm$1.24 5.26$\pm$0.71 3.22$\pm$0.36 6.55$\pm$1.04 5.54$\pm$1.33 17.00$\pm$4.71 16.65$\pm$4.33
NLL ($\downarrow$) BLoB-Mean 0.74$\pm$0.02 0.73$\pm$0.04 0.29$\pm$0.03 0.47$\pm$0.03 0.37$\pm$0.02 0.32$\pm$0.02 0.67$\pm$0.07 0.39$\pm$0.03 1.53$\pm$0.13 1.54$\pm$0.15
 + TFB (FR) 0.61$\pm$0.00 0.49$\pm$0.02 0.23$\pm$0.03 0.43$\pm$0.02 0.32$\pm$0.01 0.29$\pm$0.02 0.61$\pm$0.04 0.35$\pm$0.02 1.36$\pm$0.11 1.51$\pm$0.10
 + TFB (C-STD) 0.69$\pm$0.01 0.59$\pm$0.05 0.25$\pm$0.01 0.46$\pm$0.02 0.34$\pm$0.01 0.31$\pm$0.01 0.68$\pm$0.05 0.37$\pm$0.02 1.49$\pm$0.13 1.45$\pm$0.13
 + TFB (Final) 0.52$\pm$0.01 0.52$\pm$0.03 0.24$\pm$0.01 0.45$\pm$0.01 0.35$\pm$0.01 0.28$\pm$0.00 0.55$\pm$0.02 0.37$\pm$0.01 1.38$\pm$0.10 1.41$\pm$0.09
@@
Table 12: TFB Performances with various LLM backbones [53, 54, 12, 26], where Accuracy (ACC) and Expected Calibration Error (ECE) are reported in percentages. The MLE training for each different backbone is conducted for 2 epochs on the concatenated dataset of six commonsense reasoning tasks, with a shared hyperparameter setting; “Fewer Epochs” represents training for 1 epoch. We use $N=10$ samples for TFB during inference and rows with shading indicate training-free Bayesianization methods that use a pre-trained LoRA as their mean. $\uparrow$” and $\downarrow$” indicate that higher and lower values are preferred, respectively. Boldface and underlining denote the best and the second-best performance, respectively.
@@
Table 13: Performance of TFB when applied to variants of LoRAs [23, 67, 35, 41], where Accuracy (ACC) and Expected Calibration Error (ECE) are reported in percentages. The MLE training for each LoRA variant is conducted with pre-trained Llama3.1-8B model for 2 epochs on the concatenated dataset of six commonsense reasoning tasks, with a shared hyperparameter setting. We set the number of samples to $N=10$ for TFB during inference and rows with shading indicate training-free Bayesianization methods that use a pre-trained LoRA as their mean. $\uparrow$” and $\downarrow$” indicate that higher and lower values are preferred, respectively. Boldface denotes the best performance.
@@
Table 14: Performance of Last-Layer TFB (LL TFB) applied to LoRA on Llama3.1-8B pre-trained weights, where Accuracy (ACC) and Expected Calibration Error (ECE) are reported in percentages. The evaluation is done across six common-sense reasoning tasks with a shared hyper-parameter setting after 5 epochs. We sample $N$ times during inference in the sampling-based methods. Rows with shading indicate training-free Bayesianization methods that use a pre-trained LoRA as their mean. $\uparrow$” and $\downarrow$” indicate that higher and lower values are preferred, respectively. Boldface and underlining denote the best and the second-best performance, respectively.